# Multiplicities of weakly graded families of ideals

### v2

@@ Abstract.
In this article, we extend the notion of multiplicity for weakly graded families of ideals which are bounded below linearly. In particular, we show that the limit $e_{W}(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}(R/I_{n})}{n^{d}}$ exists where =$\mathcal{I}=\{I_{n}\}$ is a bounded below linearly weakly graded families of ideals in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$ with $\dim(N(\hat{R}))<d$. Furthermore, we prove that “volume=multiplicity” formula and Minkowski inequality hold for such families of ideals. We explore some properties of $e_{W}(\mathcal{J})$ for weakly graded families of ideals of the form $\mathcal{J}=\{(I_{n}:K)\}$ where $\{I_{n}\}$ is an $\mathfrak{m}$-primary graded family of ideals. We provide a necessary and sufficient condition for the equality in Minkowski inequality for the weakly graded family of ideals of the form =$=\{(I_{n}:K)\}$ where $\{I_{n}\}$ is a bounded filtration. Moreover, we generalize a result of Rees characterizing the inclusion of ideals with the same multiplicities for the above families of ideals. Finally, we investigate the asymptotic behaviour of the length function $\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))$ where $\{I_{n}\}$ is a filtration of ideals (not necessarily $\mathfrak{m}$-primary).
@@
Some easy examples show that the limit $e(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}({R}/{I_{n}})}{n^{d}}$ can be an irrational number for a non-Noetherian filtration =$\mathcal{I}=\{I_{n}\}$ of $\mathfrak{m}$-primary ideals [6]. There are examples of graded families of $\mathfrak{m}$-primary ideals for which the above limit does not exist in a Noetherian local ring [5]. The question of whether such limits exist has been considered by several mathematicians (see Ein, Lazarsfeld and Smith [13], Mustaţă [20]). If $R$ is a local domain which is essentially of finite type over an algebraically closed residue field then Lazarsfeld and Mustaţă [18] proved that the above limit exists for graded families of $\mathfrak{m}$-primary ideals using a method introduced by Okounkov [22]. In [5], Cutkosky proved that in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$, $e(\mathcal{I})$ exists for any graded family $\mathcal{I}$ of $\mathfrak{m}$-primary ideals if and only if $\dim N(\hat{R})<\dim R$ where $\hat{R}$ is the $\mathfrak{m}$-adic completion of $R$. He also showed that the “volume=multiplicity” formula holds for graded families of $\mathfrak{m}$-primary ideals [5]. This leads naturally to the question of whether the limit $e_{W}(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}({R}/{I_{n}})}{n^{d}}$ exists for bounded below linearly weakly graded family of ideals =$\mathcal{I}=\{I_{n}\}$ (see Definition 2.4). The main objective of this article is to prove the existence of the above limit and investigate some of the classical results about multiplicities for a bounded below linearly weakly graded family of ideals, namely Minkowski inequality and Rees’ characterization of equality of multiplicities.
@@
Minkowski inequality of multiplicities of ideals was originated in Teissier’s work on equisingularity. In his Cargèse paper [30], he conjectured that the Minkowski inequality $e(IJ)^{1/d}\leq e(I)^{1/d}+e(J)^{1/d}$ holds for any two $\mathfrak{m}$-primary ideals $I$ and $J$ in $R$ with $d\geq 1$ and proved it when $R$ is reduced Cohen-Macaulay and contains $\mathbb{Q}$ [31]. Later Rees and Sharp proved it in full generality [26]. In [6], Cutksoky, Srinivasan and the author proved that the Minkowski inequality holds for filtrations (not necessarily Noetherian) of $\mathfrak{m}$-primary ideals. Minkowski equality is further explored in [7] and [9].
@@
Rees used multiplicities of $\mathfrak{m}$-primary ideals to investigate the numerical characterization of reductions. He showed that if $R$ is a formally equidimensional local ring and $J\subset I$ are $\mathfrak{m}$-primary ideals in $R$ then the integral closures of the Rees algebras $R[It]$ and $R[Jt]$ in the polynomial ring $R[t]$ are same if and only if $e(I)=e(J)$ [25]. In [6], Cutksoky, Srinivasan and the author considered the equality of $e(\mathcal{I})$ and $e(\mathcal{J})$ for filtrations $\mathcal{I}$ and $\mathcal{J}$ (not necessarily Noetherian) of $\mathfrak{m}$-primary ideals and generalized the result of Rees [25]. Therefore it is natural to analyse the Minkowski inequality and explore the equality of multiplicities of bounded below linearly weakly graded families of ideals. In this direction, in section $3$, we show the following.
@@ 2. $(ii)$
We conclude this section by showing that for weakly graded family (not necessarily bounded below linearly) of ideals, the limit $\lim\limits_{n\to\infty}\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists and bounded by the epsilon multiplicity of the filtration $\{I_{n}\}$ under some extra assumptions on $\{I_{n}\}$.
@@
    Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$ and $K$ be an $\mathfrak{m}$-primary ideal in $R$ with $I_{1}\subset K$. Then
@@
    In particular, if $I$ is an ideal in $R$ with $\operatorname{grade}(I)\geq 1$ and $K$ is an $\mathfrak{m}$-primary ideal in $R$ with $I\subset K$ then
@@
    Suppose $\{J_{n}\}$ be a (bounded below linearly) graded family of ideals in $R$ such that $J_{1}\cap R^{o}\neq\emptyset$. Let $a\in{\mathbb{R}}_{>0}$ and $c\in J_{1}\cap R^{o}$. Then for all $m,n\geq 1$, we have
@@
    Hence $\{I_{n}=J_{\lfloor{an}\rfloor}\}$ is a (bounded below linearly) weakly graded family of ideals in $R$.
@@
    Let $A=\{K(1),\ldots,K(m)\}$ be a collection of ideals such that $K(i)\cap R^{o}\neq\emptyset$ for all $1\leq i\leq m$. Then =$\mathcal{I}=\{I_{n}=(J_{n}:K_{n})\}$ is a (bounded below linearly) weakly graded family of ideals where $K_{n}$ varies in $A$. Let $c_{i}\in K(i)\cap R^{o}$ for all $1\leq i\leq m$ and $d=c_{1}\cdots c_{m}$. Then $cd^{2}I_{n}I_{m}\subset I_{n+m}$ for all $m,n\geq 1$.
@@
We will denote the integral closure of an ideal $I$ in $R$ by $\overline{I}$. Let =$\mathcal{I}=\{I_{n}\}$ be a graded family of ideals in $R$. We say =$\mathcal{I}=\{I_{n}\}$ is a Noetherian graded family if the graded $R$-algebra $R[\mathcal{I}]=\bigoplus_{n\in{\mathbb{N}}}I_{n}t^{n}$ is a finitely generated $R$-algebra. Otherwise, we say =$\mathcal{I}=\{I_{n}\}$ is non-Noetherian. Let $\overline{R[\mathcal{I}]}$ denote the integral closure of $R[\mathcal{I}]$ in the polynomial ring $R[t]$. It is shown in [9, Lemma 3.6] that for a filtration =$\mathcal{I}=\{I_{n}\}$ the integral closure of $R[\mathcal{I}]$ in $R[t]$ is $\overline{R[\mathcal{I}]}=\bigoplus_{m\geq 0}J_{m}t^{m}$ where $\{J_{m}\}$ is the filtration
@@
A discrete valued filtration of $R$ is a filtration =$\mathcal{I}=\{I_{n}\}$ such that there exist discrete valuations $\nu_{1},\ldots,\nu_{r}$ and $a_{1},\ldots,a_{r}\in{\mathbb{R}}_{>0}$ such that for all $m\in{\mathbb{N}}$,
$I_{m}=I(\nu_{1})_{\lceil ma_{1}\rceil}\cap\cdots\cap I(\nu_{r})_{\lceil ma_{r}\rceil}.$
@@
$\{I_{m}=I(\nu_{1})_{\lceil ma_{1}\rceil}\cap\cdots\cap I(\nu_{r})_{\lceil ma_{r}\rceil}\}$
@@
In [5, Corollary 6.3], it is shown that if $R$ is analytically unramified then the epsilon multiplicity of $I$ is a limit. Epsilon multiplicity of a filtration is introduced in [10]. Epsilon multiplicity of a filtration satisfying $A(r)$ for some $r\in{\mathbb{Z}}_{>0}$ is a limit. For more details about $A(r)$ condition, see [10].
@@
In this section, we show the existence of the limit $e_{W}(\mathcal{I})=\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{n})/n^{d}$ for a bounded below linearly weakly graded family of ideals =$\mathcal{I}=\{I_{n}\}$ in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$ with $\dim(N(\hat{R}))<d$. We generalize the “multiplicity=volume” formula for a bounded below linearly weakly graded family of ideals. We also prove Minkowski inequality and show that this inequality can be strict in general. We provide a sufficient condition for the equality of the multiplicities of two bounded below linearly weakly graded families of ideals. The following lemma is well-known. For the sake of completeness, we include the proof here.
@@ 3. $(c)$
$(iv)$ Since $I_{n}\subset J_{n}$ and $\overline{I_{n}}=\overline{J_{n}}$ for all $n\gg 0$, we have $e(I_{n})=e(J_{n})$ for all $n\gg 0$ by [29, Proposition 11.2.1]. Thus we get the required result. ∎
@@
    Part $(i)$, $(ii)$ and $(iii)$ of Theorem 3.2 give alternative proofs for Theorems $10.11$, $10.14$ and $10.16$ in [12].
@@
$x\sqrt{x+1}=a_{1}x+a_{2}x^{2}+a_{3}x^{3}+\cdots\mbox{ where }a_{n+1}=\frac{(-1)^{n-1}(2n-3)!}{2^{2n-2}n!(n-2)!}.$
@@
in $R$. Let $\{I_{n}=G_{\lfloor\frac{n}{2}\rfloor}\}$ and $\{J_{n}=H_{\lfloor\frac{n}{2}\rfloor}\}$. Then $\{I_{n}\}$ and $\{J_{n}\}$ are weakly graded families of ideals and
@@
For a weakly graded family of ideals (not necessarily bounded below linearly) of the form $\{(I_{n}:K)\}$ where =$\mathcal{I}=\{I_{n}\}$ is a filtration and satisfies $A(r)$ condition for some $r\in{\mathbb{Z}}_{>0}$, we show that the limit $\lim\limits_{n\to\infty}\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists and it is bounded above by $\varepsilon(\mathcal{I})$.
@@ 3. $(iii)$
Therefore it is enough to show that $e(\mathcal{I})=e(\mathcal{I}S)\leq\lim\limits_{n\to\infty}d!\ell_{S}(S/(I_{n}S:KS))/n^{d}$. Thus we assume $R$ is a reduced complete local ring. Note that $I_{n}$ are $\mathfrak{m}$-primary ideals and hence $\operatorname{grade}(I_{n})\geq 1$ for all $n\geq 1$. Since $\mathcal{I}$ is a Noetherian filtration, by [1, Proposition 3, page 159], there exists an integer $e\geq 1$ such that $I_{en}=I_{e}^{n}$ for all $n\geq 1$. Now $K$ is an $\mathfrak{m}$-primary ideal implies there exists $t\in{\mathbb{Z}}_{>0}$ such that $I_{et}=I_{e}^{t}\subseteq K$. By [23, Theorem 4.1], for all $n\gg 0$, we have $(I_{e}^{n}:K)\subset(I_{e}^{n}:I_{e}^{t})=I_{e}^{n-t}.$ Hence using $(i)$, we get
@@
$\geq$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{e}^{n-t})/{(en)}^{d}$
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{e(n-t)})/{(en)}^{d}=e(\mathcal{I}).$
Since $R$ is an analytically unramified local ring, by [24], $\{\overline{I^{n}}\}$ is a Noetherian filtration. It is well known that $e(I)=e(\mathcal{I})$. Thus we get the required result.
@@
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{wn}:K))/{(wn)}^{d}\geq\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{w(n-1)})/{(wn)}^{d}=e(\mathcal{I}).$

### v3

@@ Abstract.
In this article, we extend the notion of multiplicity for weakly graded families of ideals which are bounded below linearly. In particular, we show that the limit $e_{W}(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}(R/I_{n})}{n^{d}}$ exists where =$\mathcal{I}=\{I_{n}\}$ is a bounded below linearly weakly graded family of ideals in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$ with $\dim(N(\hat{R}))<d$. Furthermore, we prove that “volume=multiplicity” formula and Minkowski inequality hold for such families of ideals. We explore some properties of $e_{W}(\mathcal{J})$ for weakly graded families of ideals of the form $\mathcal{J}=\{(I_{n}:K)\}$ where $\{I_{n}\}$ is an $\mathfrak{m}$-primary graded family of ideals. We provide a necessary and sufficient condition for the equality in Minkowski inequality for the weakly graded families of ideals of the form $\{(I_{n}:K)\}$ where $\{I_{n}\}$ is a bounded filtration. Moreover, we generalize a result of Rees characterizing the inclusion of ideals with the same multiplicities for the above families of ideals. Finally, we investigate the asymptotic behaviour of the length function $\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))$ where $\{I_{n}\}$ is a filtration of ideals (not necessarily $\mathfrak{m}$-primary).
@@
Some easy examples show that the limit $e(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}({R}/{I_{n}})}{n^{d}}$ can be an irrational number for a non-Noetherian filtration =$\mathcal{I}=\{I_{n}\}$ of $\mathfrak{m}$-primary ideals [6]. There are examples of graded families of $\mathfrak{m}$-primary ideals for which the above limit does not exist in a Noetherian local ring [5]. The question of whether such limits exist has been considered by several mathematicians (see Ein, Lazarsfeld and Smith [13], Mustaţă [20]). If $(R,\mathfrak{m})$ is a local domain which is essentially of finite type over an algebraically closed field $k$ with $R/\mathfrak{m}=k$ then Lazarsfeld and Mustaţă [18] proved that the above limit exists for any graded family of $\mathfrak{m}$-primary ideals using a method introduced by Okounkov [22]. In [5], Cutkosky proved that in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$, $e(\mathcal{I})$ exists for any graded family $\mathcal{I}$ of $\mathfrak{m}$-primary ideals if and only if $\dim N(\hat{R})<\dim R$ where $\hat{R}$ is the $\mathfrak{m}$-adic completion of $R$. He also showed that the “volume=multiplicity” formula holds for graded families of $\mathfrak{m}$-primary ideals [5]. This leads naturally to the question of whether the limit $e_{W}(\mathcal{I}):=\lim\limits_{n\to\infty}d!\frac{\ell_{R}({R}/{I_{n}})}{n^{d}}$ exists for a bounded below linearly weakly graded family of ideals =$\mathcal{I}=\{I_{n}\}$ (see Definition 2.4). The main objective of this article is to prove the existence of the above limit and investigate some of the classical results about multiplicities for a bounded below linearly weakly graded family of ideals, namely Minkowski inequality and Rees’ characterization of equality of multiplicities.
@@
Minkowski inequality of multiplicities of ideals was originated in Teissier’s work on equisingularity. In his Cargèse paper [30], he conjectured that the Minkowski inequality $e(IJ)^{1/d}\leq e(I)^{1/d}+e(J)^{1/d}$ holds for any two $\mathfrak{m}$-primary ideals $I$ and $J$ in $R$ with $d\geq 1$ and proved it when $R$ is reduced Cohen-Macaulay and contains $\mathbb{Q}$ [31]. Later Rees and Sharp proved it in full generality [26]. In [6], Cutkosky, Srinivasan and the author proved that the Minkowski inequality holds for filtrations (not necessarily Noetherian) of $\mathfrak{m}$-primary ideals. Minkowski equality is further explored in [7] and [9].
@@
Rees used multiplicities of $\mathfrak{m}$-primary ideals to investigate the numerical characterization of reductions. He showed that if $R$ is a formally equidimensional local ring and $J\subset I$ are $\mathfrak{m}$-primary ideals in $R$ then the integral closures of the Rees algebras $R[It]$ and $R[Jt]$ in the polynomial ring $R[t]$ are same if and only if $e(I)=e(J)$ [25]. In [6], Cutkosky, Srinivasan and the author considered the equality of $e(\mathcal{I})$ and $e(\mathcal{J})$ for filtrations $\mathcal{I}$ and $\mathcal{J}$ (not necessarily Noetherian) of $\mathfrak{m}$-primary ideals and generalized the result of Rees [25]. Therefore it is natural to analyse the Minkowski inequality and explore the equality of multiplicities of bounded below linearly weakly graded families of ideals. In this direction, in section $3$, we show the following.
@@ 2. $(ii)$
We conclude this section by showing that for a weakly graded family (not necessarily bounded below linearly) of ideals of the form $\{(I_{n}:K)\}$, the limit $\lim\limits_{n\to\infty}\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists and bounded above by the epsilon multiplicity of the filtration $\{I_{n}\}$ under some extra assumptions on $\{I_{n}\}$ and $K$.
@@
    Let =$\mathcal{I}=\{I_{n}\}$ be a Noetherian filtration of ideals in $R$ with $\operatorname{grade}(I_{1})\geq 1$ and $K$ be an $\mathfrak{m}$-primary ideal in $R$. Then
@@
    In particular, if $I$ is an ideal in $R$ with $\operatorname{grade}(I)\geq 1$ and $K$ is an $\mathfrak{m}$-primary ideal in $R$ then
@@
    Suppose $\{J_{n}\}$ is a (bounded below linearly) graded family of ideals in $R$ such that $J_{1}\cap R^{o}\neq\emptyset$. Let $a\in{\mathbb{R}}_{>0}$ and $c\in J_{1}\cap R^{o}$. Then for all $m,n\geq 1$, we have
@@
    (If $\mathfrak{m}^{sn}\subset J_{n}$ then $\mathfrak{m}^{s(\lfloor{a}\rfloor+1)n}\subset J_{\lfloor{an}\rfloor}$ for all $n\geq 1$.) Hence $\{I_{n}=J_{\lfloor{an}\rfloor}\}$ is a (bounded below linearly) weakly graded family of ideals in $R$.
@@
    Let $A=\{K(1),\ldots,K(r)\}$ be a collection of ideals in $R$ such that $K(i)\cap R^{o}\neq\emptyset$ for all $1\leq i\leq r$. Then =$\mathcal{I}=\{I_{n}=(J_{n}:K_{n})\}$ is a (bounded below linearly) weakly graded family of ideals where $K_{n}$ varies in $A$. Let $c_{i}\in K(i)\cap R^{o}$ for all $1\leq i\leq r$ and $d=c_{1}\cdots c_{r}$. Then $cd^{2}I_{n}I_{m}\subset I_{n+m}$ for all $m,n\geq 1$.
@@
We denote the integral closure of an ideal $I$ in $R$ by $\overline{I}$. Let =$\mathcal{I}=\{I_{n}\}$ be a graded family of ideals in $R$. We say =$\mathcal{I}=\{I_{n}\}$ is a Noetherian graded family if the graded $R$-algebra $R[\mathcal{I}]=\bigoplus_{n\in{\mathbb{N}}}I_{n}t^{n}$ is a finitely generated $R$-algebra. Otherwise, we say =$\mathcal{I}=\{I_{n}\}$ is non-Noetherian. Let $\overline{R[\mathcal{I}]}$ denote the integral closure of $R[\mathcal{I}]$ in the polynomial ring $R[t]$. It is shown in [9, Lemma 3.6] that for a filtration =$\mathcal{I}=\{I_{n}\}$, the integral closure of $R[\mathcal{I}]$ in $R[t]$ is $\overline{R[\mathcal{I}]}=\bigoplus_{m\geq 0}J_{m}t^{m}$ where $\{J_{m}\}$ is the filtration
@@
A discrete valued filtration of $R$ is a filtration =$\mathcal{I}=\{I_{n}\}$ such that there exist discrete valuations $\nu_{1},\ldots,\nu_{r}$ of $K$ and $a_{1},\ldots,a_{r}\in{\mathbb{R}}_{>0}$ such that $R\subset\mathcal{O}_{{\nu}_{i}}$ for all $1\leq i\leq r$ and for all $n\in{\mathbb{N}}$,
$I_{n}=I(\nu_{1})_{\lceil na_{1}\rceil}\cap\cdots\cap I(\nu_{r})_{\lceil na_{r}\rceil}.$
@@
$\{I_{n}=I(\nu_{1})_{\lceil na_{1}\rceil}\cap\cdots\cap I(\nu_{r})_{\lceil na_{r}\rceil}\}$
@@
In [5, Corollary 6.3], it is shown that if $R$ is analytically unramified then the epsilon multiplicity of $I$ is a limit. Epsilon multiplicity of a filtration is introduced in [10]. Epsilon multiplicity of a filtration =$\mathcal{I}=\{I_{n}\}$ satisfying $A(r)$ for some $r\in{\mathbb{Z}}_{>0}$ is a limit, i.e., $\varepsilon(\mathcal{I})=\lim\limits_{n\to\infty}d!\frac{\ell_{R}(H^{0}_{\mathfrak{m}}(R/I_{n}))}{n^{d}}$ exists. For more details about $A(r)$ condition, see [10].
@@
In this section, we show the existence of the limit $e_{W}(\mathcal{I})=\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{n})/n^{d}$ for a bounded below linearly weakly graded family of ideals =$\mathcal{I}=\{I_{n}\}$ in a Noetherian local ring $(R,\mathfrak{m})$ of dimension $d\geq 1$ with $\dim(N(\hat{R}))<d$. We generalize the “volume=multiplicity” formula for a bounded below linearly weakly graded family of ideals. We also prove Minkowski inequality and show that this inequality can be strict in general. We provide a sufficient condition for the equality of the multiplicities of two bounded below linearly weakly graded families of ideals. The following lemma is well-known. For the sake of completeness, we include the proof here.
@@ 3. $(c)$
$(iv)$ Since $I_{n}\subset J_{n}$ and $\overline{I_{n}}=\overline{J_{n}}$ for all $n\gg 0$, we have $e(I_{n})=e(J_{n})$ for all $n\gg 0$ by [29, Proposition 11.2.1]. Thus we get the required result by part $(ii)$. ∎
@@
    Part $(i)$, part $(ii)$ and part $(iii)$ of Theorem 3.2 give alternative proofs for Theorems $10.11$, $10.14$ and $10.16$ in [12].
@@
$x\sqrt{x+1}=a_{1}x+a_{2}x^{2}+a_{3}x^{3}+\cdots\mbox{ where }a_{n+1}=\frac{(-1)^{n-1}(2n-3)!}{2^{2n-2}n!(n-2)!}\mbox{ for }n\geq 2.$
@@
in $R$. Let $\{I_{n}=G_{\lfloor\frac{n}{2}\rfloor}\}$ and $\{J_{n}=H_{\lfloor\frac{n}{2}\rfloor}\}$. Then $\{I_{n}\}$ and $\{J_{n}\}$ are bounded below linearly weakly graded families of ideals and
@@
For a weakly graded family of ideals (not necessarily bounded below linearly) of the form $\{(I_{n}:K)\}$ where =$\mathcal{I}=\{I_{n}\}$ is a filtration and satisfies $A(r)$ condition for some $r\in{\mathbb{Z}}_{>0}$, we show that the limit $\lim\limits_{n\to\infty}\ell_{R}(H_{\mathfrak{m}}^{0}(R/(I_{n}:K)))/{n}^{d}$ exists and it is bounded above by $\varepsilon(\mathcal{I})$ if $K$ is $\mathfrak{m}$-primary.
Remark 4.1.
Suppose =$\mathcal{I}=\{I_{n}\}$ is a Noetherian filtration of ideals in a Noetherian local ring $(R,\mathfrak{m})$. Then by [1, Proposition 3, page 159], there exists an integer $m\geq 1$ such that $I_{mn}=I_{m}^{n}$ for all $n\geq 1$. Let $K$ be an $\mathfrak{m}$-primary ideal in $R$ and $I_{1}\neq R$. Then there exists an integer $r\geq 1$ such that $I_{mr}=I_{m}^{r}\subset\mathfrak{m}^{r}\subset K$. Let $e=mr$. Then for all $n\geq 1$, we have $I_{en}=I_{mrn}=I_{m}^{rn}=I_{mr}^{n}=I_{e}^{n}$.
@@ 3. $(iii)$
Therefore it is enough to show that $e(\mathcal{I})=e(\mathcal{I}S)\leq\lim\limits_{n\to\infty}d!\ell_{S}(S/(I_{n}S:KS))/n^{d}$. Thus we assume $R$ is a reduced complete local ring. Note that $I_{n}$ are $\mathfrak{m}$-primary ideals and hence $\operatorname{grade}(I_{n})\geq 1$ for all $n\geq 1$. Since $\mathcal{I}$ is a Noetherian filtration, by Remark 4.1, there exists an integer $e\geq 1$ such that $I_{e}\subset K$ and $I_{en}=I_{e}^{n}$ for all $n\geq 1$. By [23, Theorem 4.1], for all $n\gg 0$, we have $(I_{e}^{n}:K)\subset(I_{e}^{n}:I_{e})=I_{e}^{n-1}.$ Hence using $(i)$, we get
@@
$\geq$ $\lim\limits_{n\to\infty}\big{[}\Big{(}d!\ell_{R}(R/I_{e}^{n-1})/{(e(n-1))}^{d}\Big{)}\Big{(}{(e(n-1))}^{d}/{(en)}^{d}\Big{)}\big{]}$
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{e(n-1)})/{(e(n-1))}^{d}=e(\mathcal{I}).$
Since $R$ is an analytically unramified local ring, by [24], =$\mathcal{I}=\{\overline{I^{n}}\}$ is a Noetherian filtration. It is well known that $e(I)=e(\mathcal{I})$. Thus we get the required result.
@@
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/(I_{wn}:K))/{(wn)}^{d}$
$\geq$ $\lim\limits_{n\to\infty}\big{[}\Big{(}d!\ell_{R}(R/I_{w(n-1)})/{(w(n-1))}^{d}\Big{)}\Big{(}{(w(n-1))}^{d}/{(wn)}^{d}\Big{)}\big{]}$
$=$ $\lim\limits_{n\to\infty}d!\ell_{R}(R/I_{w(n-1)})/{(w(n-1))}^{d}=e(\mathcal{I}).$